Convert 1.43 x 10^24 atoms of carbon to moles of carbon.

1.43 x 10^24 atoms= ________ mol C
1
Expert's answer
2020-10-02T14:20:08-0400


We know,

1 moles of C = Avogadro number(NA) of carbon atom


=6.023×10^23 atoms of carbon


Hence,


6.023×10^23 atoms of carbon =1 mole C


1 atoms of carbon = 1/(6.023×10^23) mole C


1.43 x 10^24 atoms of carbon =(1.43 x 10^24 )/(6.023×10^23) moles C

= 2.274 mol. C

(Upto three significant decimal place)


Hence,

1.43 x 10^24 atoms= 2.274 mol C